Find the mean and standard deviation of the sample data set. Show your complete solution.

A. Ages of Shoppers. The ages (in years) of a random sample of shoppers at a clothing outlet.

16 18 19 17 14 15 17 17 17 16 19 22 24 14 16 14 17 16 14 18


1
Expert's answer
2020-12-15T14:24:44-0500

Mean

"\bar x=\frac{\sum_i^n x_i}{n}"

="\frac{16+18+19+...+18}{20}"

="\frac{340}{20}"

=17

standard deviation

"s=\sqrt{\frac{\sum_i^n (x_i-\bar x)^2}{n-1}}"

="\sqrt{\frac{(16-17)^2+(18-17)^2+(19-17)^2+...+(18-17)^2}{20-1}}"

="\sqrt{\frac{128}{19}}"

=2.5955